OpenStudy (anonymous):

find a power series representation for the function and determine the interval of convergence \[f \left( x \right)=\left( x \right)\div \left( 2x ^{2}+1 \right)\]

OpenStudy (somethingawesome):

We're going to need to use the geometric series expansion, that is, Sum x^n = 1/(1-x). What this means is we're going to have to play with f(x) until it looks like 1/(1-something). Why not make it f(x) = x/(1- (-2x^2))? Then \[\frac{1}{1-(-2x^2)} = \sum_{n=0}^\infty (-2x^2)^n\] \[= \sum_{n=0}^\infty (-1)^n 2^n x^{2n}\] But this isn't quite f(x), we want an x on top, so multiply both sides by x: \[f(x) = \frac{x}{1-(-2x^2)} = \sum_{n=0}^\infty (-1)^n 2^n x^{2n+1}.\] For the interval of convergence, use the ratio test on this series. Can you take it from here?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

\[\sum_{0}^{\infty}\left( 2^{n +1}x ^{2n +2} \right)\div \left( 2^{n}x ^{2n +1} \right)\] is this right so far?

OpenStudy (somethingawesome):

Yep, except that it should be x^{2(n+1)+1} or x^{2n+3} on top.

OpenStudy (anonymous):

ok im always confused about this next step, do we plug in infinity?

OpenStudy (somethingawesome):

Always simplify as much as you can before taking any limits. So we have \[\lim_{n\rightarrow \infty} \frac{2^{n+1} x^{2n+3}}{2^n x^{2n+1}}\] that we can cancel some terms from.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}2^{n}x ^{2n +2}\] ?

OpenStudy (somethingawesome):

We subtract exponents on like terms, so 2^{n+1}/2^n = 2^{n+1 - n} = 2, and x^{2n+3}/x^{2n+1} = x^2. We should have \[\lim_{n\rightarrow\infty} \left| 2 x^2 \right|\] There are no more n's, so taking the limit doesn't do much. Set this less than 1 and solve for |x| for the radius of convergence.

OpenStudy (anonymous):

thank you so much. im having so much trouble with this subject its driving me insane

OpenStudy (somethingawesome):

Keep at it, study and practice, and you'll do fine! You've made it this far, after all.